# Offline Robotic World Model: Learning Robotic Policies without a Physics Simulator

### v1

@@ 4.2 Policy Optimization on Uncertainty-Penalized Imagination
To ensure robust policy learning in offline MBRL, we introduce Model-Based Offline Policy Optimization with Proximal Policy Optimization (MOPO-PPO), an adaptation of the MOPO framework [15] to PPO [42]. MOPO-PPO allows cautious exploration beyond the behavioral distribution while penalizing transitions with high epistemic uncertainty. Specifically, given the uncertainty estimator $u_{p_{\phi}}$ from Eq. 4, we modify the reward function to discourage the policy from exploiting unreliable model predictions:
@@
The training process of MOPO-PPO using RWM-O is outlined in Fig. 1 and Algorithm 1.
Algorithm 1 Offline policy optimization with RWM-O
@@
4:  for learning iterations $=1,2,\dots$ do
@@
By integrating uncertainty-aware dynamics modeling with policy optimization, RWM-O and MOPO-PPO enable safer and more reliable offline MBRL. Explicit uncertainty estimation mitigates distribution shift, allowing cautious generalization beyond the dataset while ensuring stability through trust-region optimization. By leveraging uncertainty-aware rollouts, MOPO-PPO enhances policy robustness and efficiency, advancing scalable offline RL for real-world robotics.
@@ 5 Experiments
We evaluate RWM-O and MOPO-PPO through a comprehensive set of experiments spanning diverse robotic platforms, tasks, and dataset sources. Our objective is to assess the capability of RWM-O in quantifying uncertainty during long-horizon rollouts and to demonstrate the effectiveness of MOPO-PPO in improving policy performance through uncertainty-aware optimization.
@@ 5.1 Autoregressive Uncertainty Estimation
A reliable estimate of epistemic uncertainty is critical for ensuring robust policy learning in model-based offline RL. In particular, the ability of a world model to quantify its own uncertainty over long-horizon rollouts determines the trustworthiness of its synthetic experience. To evaluate this property in RWM-O, we analyze the alignment between its predicted epistemic uncertainty and actual model prediction errors during autoregressive trajectory forecasting on ANYmal D. The observation and action spaces of the world model are detailed in Table S1 and Table S3.
We train the world model on offline data collected from a velocity-tracking policy, operating at a control frequency of $50\,Hz$. The model is trained with a history horizon of $M=32$ and a prediction horizon of $N=8$, using an ensemble of five networks to estimate epistemic uncertainty. To evaluate uncertainty estimation, we compute epistemic uncertainty as the variance of the ensemble predictions, following Eq. 4. Aleatoric uncertainty is estimated as the mean of the predicted standard deviations across ensemble members. Details of the network architecture and training setup are provided in Sec. A.2.1 and Sec. A.3.1. The predictive error, epistemic uncertainty, and aleatoric uncertainty over autoregressive rollouts are visualized in Fig. 2.
Figure 2: Autoregressive trajectory prediction (left) and uncertainty estimation (right) by RWM-O. Solid lines represent ground truth trajectories, while dashed lines denote predicted state evolution. Predictions commence at $t=32$ using historical observations, with future observations predicted autoregressively by feeding prior predictions back into the model.
At $t=32$, the model transitions from conditioning on ground-truth observations to fully autoregressive rollouts, wherein predicted states are recursively fed back into the model. As expected, the accumulated prediction error (grey) increases over time due to compounding model inaccuracies. Importantly, the estimated epistemic uncertainty (dark blue) closely follows the trend of the prediction error, demonstrating that RWM-O effectively captures uncertainty in regions where the model generalization deteriorates. In contrast, the aleatoric uncertainty (light blue) remains low, reflecting small stochasticity in the environment. The strong correlation between epistemic uncertainty and model prediction error justifies its role as a trust metric for policy optimization. By penalizing high-uncertainty transitions, MOPO-PPO prevents policy exploitation in poorly modeled regions of the state space, mitigating the risk of compounding errors in long-horizon rollouts. These results validate the proposed uncertainty-aware regularization strategy, reinforcing the efficacy of RWM-O in enabling robust policy learning in offline RL.
@@ 5.2 Uncertainty-Penalized Imagination
To evaluate the role of uncertainty-aware regularization in policy learning, we train RWM-O on an offline dataset collected from an expert velocity-tracking policy on ANYmal D. We then optimize policies with MOPO-PPO using different values of the uncertainty penalty coefficient $\lambda$ in Eq. 5, which controls the trade-off between exploration and model reliability. Figure 3 shows the evolution of the imagination reward and epistemic uncertainty over training iterations, as well as the final policy evaluation performance.
@@ 5.3 Generality across Robotic Environments
To evaluate the generality and robustness of MOPO-PPO with RWM-O across diverse robotic environments, we compare its performance against widely adopted offline RL baselines, including Conservative Q-Learning (CQL) [43] with Soft Actor-Critic (SAC) [44], MBPO [27], MBPO-PPO [26], and MOPO [15]. These baselines represent state-of-the-art approaches in model-free and model-based offline RL. For a fair comparison, we train MBPO and MBPO-PPO in an offline setting, treating them as uncertainty-unaware counterparts to MOPO and MOPO-PPO.
We benchmark these methods on three robotic tasks in simulation: Reach-Franka (manipulation), Velocity-G1 (humanoid locomotion), and Velocity-ANYmal-D (quadruped locomotion), spanning a broad range of dynamics complexities and control challenges. Each method is trained on four dataset types: Random: trajectories collected from a randomly initialized policy, Medium: trajectories from a partially trained PPO policy, Expert: trajectories from a fully trained PPO policy, Mixed: a combination of replay buffer data from PPO training at various stages. All methods are trained and evaluated under identical conditions. The normalized episodic rewards for each method across different environments and dataset types are presented in Fig. 4.
@@
While CQL and SAC-based model-based methods perform competitively in the manipulation task, they demonstrate significantly weaker performance in locomotion tasks, where the complexity of dynamic interactions amplifies the limitations of value-based learning. As highlighted in Sec. 5.2, MBPO and MBPO-PPO, which lack explicit uncertainty estimation, tend to overfit to model inaccuracies, leading to overconfident policies. In manipulation tasks, this overfitting is mitigated when dataset coverage is sufficient. However, in locomotion tasks, it results in catastrophic failures, including robot collisions and falls.
@@
In contrast, MOPO-PPO consistently outperforms baselines across all tasks, particularly in Velocity-G1 and Velocity-ANYmal-D, where uncertainty-aware rollouts significantly enhance policy stability. By penalizing high-uncertainty transitions, MOPO-PPO prevents policies from exploiting hallucinated model predictions, leading to safer and more reliable locomotion behaviors. The deployment of the learned policy is visualized in Fig. S5. The strong performance of MOPO-PPO across different dataset types and environments demonstrates its effectiveness in mitigating distribution shift, reinforcing its applicability to real-world robotic control.
5.4 Offline Learning from Real-World Data
To validate the real-world applicability of RWM-O and MOPO-PPO, we train a locomotion policy entirely from offline data collected on the ANYmal D robot, without any reliance on a physics simulator. We develop an autonomous data collection pipeline with minimal human intervention in Sec. A.4.2. This experiment demonstrates that our approach eliminates the sim-to-real gap by learning directly from real-world experience. By leveraging uncertainty-aware rollouts, MOPO-PPO enables safe and effective policy optimization purely from offline data. A video demonstrating the learned policy running on hardware is available on our project webpage.
@@ 6 Conclusion
In this work, we introduce RWM-O, an uncertainty-aware MBRL framework designed to improve offline policy learning. By explicitly estimating epistemic uncertainty through an ensemble-based approach, RWM-O enables uncertainty-aware rollouts, reducing the risk of overfitting to inaccurate model predictions. We further integrate this uncertainty estimation into policy optimization with MOPO-PPO, penalizing unreliable transitions and enhancing policy stability. Our experiments demonstrate that RWM-O improves policy generalization, robustness, and safety across various robotic tasks, outperforming standard model-based and model-free offline RL baselines. Notably, our approach enables policy learning purely from real-world data, eliminating the need for high-fidelity physics simulators and addressing a key limitation of current RL methodologies. By bridging the gap between model-based offline RL and real-world deployment, RWM-O and MOPO-PPO enhance the scalability and data efficiency of RL for robotics, paving the way for more reliable and adaptable policy learning.
Limitations
While RWM-O improves offline policy learning by incorporating uncertainty estimation, several challenges remain. First, the accuracy of the learned dynamics model is inherently limited by the quality and diversity of the offline dataset. If the dataset lacks sufficient coverage of the state-action space, the model may struggle with generalization, leading to unreliable predictions in unobserved regions. Second, although our ensemble-based uncertainty estimation helps mitigate distribution shift, it does not entirely eliminate compounding errors over long-horizon rollouts. The effectiveness of the uncertainty penalty in MOPO-PPO relies on careful tuning of hyperparameters, such as the penalty weight, which may require task-specific adjustments. Finally, while RWM-O and MOPO-PPO remove the need for a physics simulator, the reliance on offline real-world data means the policy cannot explore beyond the dataset distribution. In cases where real-world data is insufficient, supplementing it with simulation data could improve the model’s coverage, but this reintroduces dependence on a physics simulator. This trade-off can limit applicability in scenarios requiring extensive exploration or rapid adaptation to new environments.
@@ Acknowledgments
This research was supported by the ETH AI Center.
@@ A.1.1 Observation and action spaces
The observation space for the ANYmal world model is composed of base linear and angular velocities $v$, $\omega$ in the robot frame, measurement of the gravity vector in the robot frame $g$, joint positions $q$, velocities $\dot{q}$ and torques $\tau$ as in Table S1.
@@
Entry Symbol Dimensions
base linear velocity $v$ 0:3
base angular velocity $\omega$ 3:6
projected gravity $g$ 6:9
joint positions $q$ 9:21
joint velocities $\dot{q}$ 21:33
joint torques $\tau$ 33:45
The privileged information is used to provide an additional learning objective that implicitly embeds critical information for accurate long-term predictions. The space is composed of knee and foot contacts as in Table S2.
@@
Entry Symbol Dimensions
knee contact -- 0:4
foot contact -- 4:8
@@
Entry Symbol Dimensions
joint position targets $q^{*}$ 0:12
The observation space for the ANYmal velocity tracking policy is composed of base linear and angular velocities $v$, $\omega$ in the robot frame, measurement of the gravity vector in the robot frame $g$, velocity command $c$, joint positions $q$ and velocities $\dot{q}$ as in Table S4.
@@
Entry Symbol Dimensions
base linear velocity $v$ 0:3
base angular velocity $\omega$ 3:6
projected gravity $g$ 6:9
velocity command $c$ 9:12
joint positions $q$ 12:24
joint velocities $\dot{q}$ 24:36

### v3

@@ 4.2 Policy Optimization on Uncertainty-Penalized Imagination
To ensure robust policy learning in offline MBRL, we introduce Model-Based Offline Policy Optimization with Proximal Policy Optimization (MOPO-PPO), an adaptation of the MOPO framework (Yu et al., 2020) to PPO (Schulman et al., 2017). MOPO-PPO allows cautious exploration beyond the behavioral distribution while penalizing transitions with high epistemic uncertainty. Specifically, given the uncertainty estimator $u_{p_{\phi}}$ from Eq. 4, we modify the reward function to discourage the policy from exploiting unreliable model predictions:
@@
The training process of MOPO-PPO using RWM-U is outlined in Fig. 2 and Algorithm 1. We provide more detailed training setups in Sec. A.3.
Algorithm 1 Offline policy optimization with RWM-U
@@
4:for learning iterations $1,2,\dots$ do
@@
By integrating uncertainty-aware dynamics modeling with policy optimization, RWM-U and MOPO-PPO enable safer and more reliable offline MBRL. Explicit uncertainty estimation mitigates distribution shift, allowing cautious generalization beyond the dataset while ensuring stability through trust-region optimization. By leveraging uncertainty-aware rollouts, MOPO-PPO enhances policy robustness and efficiency, advancing scalable offline RL for real-world robotics.
@@ 5 Experiments
We evaluate RWM-U and MOPO-PPO through a comprehensive set of experiments spanning diverse simulated and real robotic platforms, tasks, and dataset sources. Our objective is to assess the capability of RWM-U in quantifying uncertainty during long-horizon rollouts and to demonstrate the effectiveness of MOPO-PPO in improving policy performance through uncertainty-aware optimization in real robotic settings. We detail our distinct challenges in Sec. A.5.
@@ 5.1 Autoregressive Uncertainty Estimation
A reliable estimate of epistemic uncertainty is critical for ensuring robust policy learning in offline MBRL. In particular, the ability of a world model to quantify its own uncertainty over long-horizon rollouts determines the trustworthiness of its synthetic experience. To evaluate this property in RWM-U, we analyze the alignment between its predicted epistemic uncertainty and actual model prediction errors during autoregressive trajectory forecasting on ANYmal D. The observation and action spaces of the world model are detailed in Table S2 and Table S4.
We train the world model on offline data collected from a velocity-tracking policy, operating at a control frequency of 50Hz50\,Hz. The model is trained with a history horizon of $M=32$ and a prediction horizon of $N=8$, using an ensemble of five networks to estimate epistemic uncertainty. To evaluate uncertainty estimation, we compute epistemic uncertainty as the variance of the ensemble predictions, following Eq. 4. Aleatoric uncertainty is estimated as the mean of the predicted standard deviations across ensemble members. Details of the network architecture and training setup are provided in Sec. A.2.1 and Sec. A.3.1. The predictive error, epistemic uncertainty, and aleatoric uncertainty over autoregressive rollouts are visualized in Fig. 3.
Figure 3: Autoregressive trajectory prediction (left) and uncertainty estimation (right) by RWM-U. Solid lines represent ground truth, while dashed lines denote predicted state evolution. Predictions commence at $t=32$ using historical observations, with future observations predicted autoregressively by feeding prior predictions back into the model. The epistemic uncertainty estimate by RWM-U aligns with the long-horizon prediction error and thus sets a reliable metric in policy training.
At $t=32$, the model transitions from conditioning on ground-truth observations to fully autoregressive rollouts, wherein predicted states are recursively fed back into the model. As expected, the accumulated prediction error (grey) increases over time due to compounding model inaccuracies. Importantly, the estimated epistemic uncertainty (dark blue) closely follows the trend of the prediction error, demonstrating that RWM-U effectively captures uncertainty in regions where the model generalization deteriorates. In contrast, the aleatoric uncertainty (light blue) remains low, reflecting small stochasticity in the environment. The strong correlation between epistemic uncertainty and model prediction error justifies its role as a trust metric for policy optimization. By penalizing high-uncertainty transitions, MOPO-PPO prevents policy exploitation in poorly modeled regions of the state space, mitigating the risk of compounding errors in long-horizon rollouts. These results validate the proposed uncertainty-aware regularization strategy, reinforcing the efficacy of RWM-U in enabling robust policy learning in offline RL.
@@ 5.2 Uncertainty-Penalized Imagination
To evaluate the role of uncertainty-aware regularization in policy learning, we train RWM-U on an offline dataset collected from velocity-tracking policies on ANYmal D and Unitree G1. We then optimize policies with MOPO-PPO using different values of the uncertainty penalty coefficient $\lambda$ in Eq. 5, which is a hyperparameter that controls the trade-off between exploration and model reliability. Figure 4 shows the evolution of the imagination reward and epistemic uncertainty over training iterations, as well as the final policy evaluation performance.
@@ 5.3 Generality across Robotic Environments
To evaluate the generality and robustness of MOPO-PPO with RWM-U across diverse robotic environments, we compare its performance against widely adopted offline RL baselines, including Conservative Q-Learning (CQL) (Kumar et al., 2020), MBPO (Janner et al., 2019), MBPO-PPO (Li et al., 2025), and MOPO (Yu et al., 2020). These baselines represent state-of-the-art approaches in model-free and model-based offline RL. For a fair comparison, we train MBPO and MBPO-PPO in an offline setting, treating them as uncertainty-unaware counterparts to MOPO and MOPO-PPO as detailed in Sec. A.2.3.
We benchmark these methods on three robotic tasks in simulation: Reach-Franka (manipulation), Velocity-ANYmal-D (quadruped locomotion) and Velocity-Unitree-G1 (humanoid locomotion), spanning a broad range of dynamics complexities and control challenges. Each method is trained on four dataset types: Random: trajectories collected from a randomly initialized policy, Medium: trajectories from a partially trained PPO policy, Expert: trajectories from a fully trained PPO policy, Mixed: a combination of replay buffer data from PPO training at various stages. All methods are trained and evaluated under identical conditions. The normalized episodic rewards for each method across different environments and dataset types are presented in Fig. 5.
@@
While CQL and SAC-based model-based methods perform competitively in the manipulation task, they demonstrate significantly weaker performance in locomotion tasks, where the complexity of dynamic interactions amplifies the limitations of value-based learning (Sec. A.5). As highlighted in Sec. 5.2, MBPO and MBPO-PPO, which lack explicit uncertainty estimation, tend to overfit to model inaccuracies, leading to overconfident policies. In manipulation tasks, this overfitting is mitigated when dataset coverage is sufficient. However, in locomotion tasks, it results in catastrophic failures, including robot collisions and falls.
@@
In contrast, MOPO-PPO consistently outperforms baselines across all tasks, particularly in Velocity-ANYmal-D and Velocity-Unitree-G1, where uncertainty-aware rollouts significantly enhance policy stability. By penalizing high-uncertainty transitions, MOPO-PPO prevents policies from exploiting hallucinated model predictions, leading to safer and more reliable locomotion behaviors. The deployment of the learned policy is visualized in Fig. 6 with supplementary videos available on our webpage. The strong performance of MOPO-PPO across different dataset types and environments demonstrates its effectiveness in mitigating distribution shift, reinforcing its applicability to real-world robotic control.
Figure 6: MOPO-PPO policy deployment learned from offline Mixed dataset across various environments. Coordinate and arrow markers denote the end-effector pose and base velocity, respectively.
5.4 Offline Learning with Real-World Data
To validate the real-world applicability of RWM-U and MOPO-PPO, we train quadruped and humanoid locomotion policies entirely from offline datasets that combine different proportions of simulation and real-world data. The resulting policy deployment is depicted in Fig. 7.
Figure 7: MOPO-PPO policy deployment on ANYmal D and Unitree G1 under varying levels of model-uncertainty penalization. The green arrow indicates the commanded velocity target. With insufficient uncertainty penalization, the robots fail to maintain stable locomotion, leading to unintended collisions or complete failure to walk. With excessive penalization, the policies become overly conservative, causing the robot to remain stationary or execute only small, ineffective motions. With appropriate uncertainty penalization, the robots explore novel behaviors in imagination while staying within reliable regions of the learned model, resulting in high-performance policies comparable to state-of-the-art sim-to-real methods.
For simulation, we use the Mixed dataset collected from training a velocity-tracking policy under heavy randomization. For real-world data, we deploy the same policy and collect trajectories using an autonomous pipeline with minimal human intervention, as described in Sec. A.4. For comparison with the state-of-the-art sim-to-real pipeline, we also train an online model-free policy in a carefully tuned simulator.
Table 1 reports the normalized real-world policy performance on ANYmal D, where the data-collecting policy achieves an average score of $0.79$. Remarkably, MOPO-PPO surpasses this baseline even when trained exclusively on simulated transitions (first column), highlighting the ability of offline MBRL to distill stronger policies than those present in the dataset. Incorporating a moderate amount of real-world data further improves performance by reducing domain shift, yielding the best result of $0.91\pm 0.03$ with 800800K simulated and 200200K real transitions, surpassing the policies trained with online model-free methods purely in simulation. However, excessive reliance on real-world data reduces performance and increases failure rates, reflecting the limited diversity of real trajectories that lack informative but risky behaviors (e.g., failure modes) necessary for robust world model and policy learning.
Table 1: Normalized policy performance on ANYmal D with different data mixtures
Sim 1M 800K 600K 400K collecting online model-
Real 0 200K 400K 600K policy free policy
Perf. 0.82 $\pm$ 0.02 0.91 $\pm$ 0.03 0.78 $\pm$ 0.02 0.59 $\pm$ 0.04 0.79 $\pm$ 0.02 0.88 $\pm$ 0.01
We additionally study the effect of uncertainty penalization during policy training in Fig. 7. With insufficient penalization ($\lambda=0.2$), the robot overfits to model errors, producing unstable gaits and frequent collisions despite high imagined rewards. Conversely, excessive penalization ($\lambda=2.0$) constrains the policy to low-uncertainty regions, resulting in overly conservative behaviors such as standing still or minimal, ineffective movements. A well-calibrated penalty ($\lambda=1.0$) strikes a balance: the imagination reward increases steadily while epistemic uncertainty remains controlled, enabling the robot to explore effectively in imagination while remaining in reliable regions of the learned dynamics model. This calibration leads to stable and robust locomotion on hardware. Supplementary videos demonstrating the learned policies on hardware are available on our project webpage.
6 Limitations
Although RWM-U and MOPO-PPO advance offline MBRL for robotics, some limitations remain. As with all offline RL approaches, policy quality is ultimately bounded by the dataset: limited state-action coverage and compounding model errors over long horizons can still degrade generalization. Our uncertainty-aware training substantially mitigates these effects but cannot eliminate them entirely. In addition, although our method enables training with offline real-world datasets, such data often lack risky but informative transitions that are essential for world model and policy learning. In practice, simulated environments remain valuable as a complementary source for generating these transitions safely. This highlights a trade-off between realism and diversity: real-world data mitigates domain shift, while simulation can enrich coverage of rare but critical experiences.
@@ 7 Conclusion
In this work, we introduce a principled pipeline for making offline MBRL effective in realistic robotic settings. Our approach, RWM-U combined with MOPO-PPO, addresses the challenges of compounding model errors and distribution shift by extending autoregressive world models with epistemic uncertainty estimation and integrating uncertainty-penalized policy optimization into a stable on-policy framework. This enables long-horizon rollouts where uncertainty is effectively propagated, allowing policies to generalize cautiously beyond the dataset while maintaining robustness to model errors. Through experiments across simulated manipulation and locomotion tasks and on real ANYmal D and Unitree G1 robots, we demonstrated that our method consistently outperforms model-free and uncertainty-unaware model-based baselines, achieving, to our knowledge, the first successful deployment of offline MBRL on physical hardware. These results establish that offline RL, when combined with principled modeling and uncertainty handling, can serve as a practical framework for reusing past robotic experience.
Ethics statement
This work does not involve human subjects or sensitive data. All experiments are conducted in simulation or on dedicated robotic hardware operated by the authors, with no use of third-party datasets. The research complies with the Code of Ethics of the venue. The proposed framework provides a robust and scalable method for learning world models tailored to complex robotic tasks. This can benefit domains such as healthcare, disaster response, and logistics, and reduce environmental and hardware costs associated with physical experimentation. Potential risks include misuse of the method in surveillance or autonomous enforcement systems, and the acceleration of automation in labor-sensitive sectors. While such uses are not intended or explored in this work, the authors acknowledge the dual-use potential of generalizable control methods. To mitigate safety risks, policy training occurs entirely in imagination, and deployment is limited to policies validated under domain shifts. Failure events are explicitly modeled and used to terminate unsafe rollouts.
Reproducibility statement
The experiment results presented in this work can be reproduced with the implementation details provided in Sec. A.1, Sec. A.2, and Sec. A.3. The code will be open-sourced on the project page upon acceptance.
@@ Acknowledgments
This research was supported by the ETH AI Center and the Swiss National Science Foundation through the National Centre of Competence in Automation (NCCR automation).
@@ A.1.1 Observation and action spaces
The observation space for the ANYmal D and Unitree G1 world model is composed of base linear and angular velocities $v$, $\omega$ in the robot frame, measurement of the gravity vector in the robot frame $g$, joint positions $q$, velocities $\dot{q}$ and torques $\tau$ as in Table S2.
@@
Entry Symbol Dimensions Entry Symbol Dimensions
ANYmal D Unitree G1
base linear velocity $v$ 0:3 base linear velocity $v$ 0:3
base angular velocity $\omega$ 3:6 base angular velocity $\omega$ 3:6
projected gravity $g$ 6:9 projected gravity $g$ 6:9
joint positions $q$ 9:21 joint positions $q$ 9:38
joint velocities $\dot{q}$ 21:33 joint velocities $\dot{q}$ 38:67
joint torques $\tau$ 33:45 joint torques $\tau$ 67:96
The privileged information is used to provide an additional learning objective that implicitly embeds critical information for accurate long-term predictions, as detailed in Table S3.
@@
Entry Symbol Dimensions Entry Symbol Dimensions
ANYmal D Unitree G1
knee contact - 0:4 body contact - 0:26
foot contact - 4:8 foot height - 26:28
foot velocity - 28:30
@@
Entry Symbol Dimensions Entry Symbol Dimensions
ANYmal D Unitree G1
joint position targets $q^{*}$ 0:12 joint position targets $q^{*}$ 0:29
The observation space for the ANYmal D and Unitree G1 velocity tracking policy is composed of base linear and angular velocities $v$, $\omega$ in the robot frame, measurement of the gravity vector in the robot frame $g$, velocity command $c$, joint positions $q$ and velocities $\dot{q}$ as in Table S5.
@@
Entry Symbol Dimensions Entry Symbol Dimensions
ANYmal D Unitree G1
base linear velocity $v$ 0:3 base linear velocity $v$ 0:3
base angular velocity $\omega$ 3:6 base angular velocity $\omega$ 3:6
projected gravity $g$ 6:9 projected gravity $g$ 6:9
velocity command $c$ 9:12 velocity command $c$ 9:12
joint positions $q$ 12:24 joint positions $q$ 12:41
joint velocities $\dot{q}$ 24:36 joint velocities $\dot{q}$ 41:70
last actions $a^{\prime}$ 36:48 last actions $a^{\prime}$ 70:99